|  August 31, 2014  |  
Fair   49.0F  |  Forecast »
Bookmark and Share Email this page Email Print this page Print Feed Feed

Repsol Nabors Blowout at Qugruk 2 UPDATE Situation Report No. 14

ALASKA DEPARTMENT OF ENVIRONMENTAL CONSERVATION
Division of Spill Prevention and Response
Prevention and Emergency Response Program

SITUATION REPORT

Repsol Q2 Pad Gas and Mud Release
SITREP # 14

SPILL NUMBER: 12399904601

TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.

TIME/DATE OF SITUATION REPORT: 3:30 PM on February 27, 2012

TYPE/AMOUNT OF PRODUCT SPILLED: Repsol estimates that approximately 42,000 gallons (1,000 barrels) of fresh-water-based drilling mud have been released to the ice drilling pad and adjacent snow-covered tundra. An unknown amount of gas has been released. The mud that was released is a mixture of minerals (primarily barium sulfate) and water. Drilling mud is used to lubricate the drill, control formation pressure, and remove cuttings from the bore. No oil has been spilled.

INCIDENT LOCATION: Qugruk #2 pad (Q2 pad), on the Colville River Delta, approximately 18 miles northeast of Nuiqsut and approximately 150 miles southeast of Barrow (70° 27’ 19” N, 150° 44’ 52” W).

CAUSE OF SPILL: Repsol contractor Nabors Drilling was drilling an exploratory well when the drill penetrated a shallow gas pocket at a depth of 2,523 feet, resulting in a gas kick. The gas kick drove drilling mud out of the well and through the gas diverter onto the ice pad and adjacent snow-covered tundra. The gas diverter is a pipe that extends approximately 75 feet from the drilling rig toward the south edge of the pad. Additional mud was pumped into the well in an attempt to control it, but that mud was also forced out by the gas.

POTENTIAL RESPONSIBLE PARTY (PRP): Repsol E&P USA

   PHOTO: ADEC

The cold start generator of the drill rig undergoing cleaning.

 

RESPONSE ACTION: Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

Technicians will inspect the cold start generator today. The cold start generator will be the first generator to be started and supplies the power needed for the eventual re-start of the rig. The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig’s power components, the time of the re-start cannot be determined at this time.

Repsol reports that 232 barrels (9,744 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,260 barrels (52,920 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Another one hundred and twenty five cubic-yards of solid waste were also manifested and shipped offsite yesterday.

SOURCE CONTROL: The well ceased flowing at about 9:00 PM on February 16. Preparations for a well-control operation are ongoing.

RESOURCES AFFECTED: Snow-covered tundra and ice pad. Drilling mud can affect tundra plants by changing soil pH and salinity. Brackish water produced by the well may also increase soil salinity.

FUTURE PLANS AND RECOMMENDATIONS: Cleanup cannot safely begin until the well is under control. The well-control team will provide cleanup personnel approximately 24-hours notice in advance of completion of the well-control operation. Plans are in place for delineating and cleaning up the spill and necessary personnel crews and equipment are on standby. A reduced incident management team will continue to monitor the operation from Anchorage and two ADEC responders will remain at Palm Pad until cleanup operations are imminent, at which time additional staff will be recalled.

UNIFIED COMMAND AND PERSONNEL:
Incident Commander: Jeremy Michels, O’Brien’s Response Management
S.O.S.C.: Tom DeRuyter, ADEC
F.O.S.C.: Matt Carr, EPA
L.O.S.C.: Gordon Brower, North Slope Borough
Field SOSC: Wes Ghormley, ADEC

WEATHER: Today: Mostly cloudy. Patchy fog. Highs 23 to 30 below. Northeast winds around 10 mph.
Tonight: Mostly cloudy. Patchy fog in the evening. Lows 35 to 40 below. West winds around
10 mph.

TIME/DATE OF THE NEXT SITUATION REPORT: 3:30 PM on February 28, 2012

FOR ADDITIONAL INFORMATION CONTACT: Ty Keltner, Public Information Officer, ADEC (907) 465-5009

Photographs and other spill information will be available for viewing at: http://www.state.ak.us/dec/spar/perp/index.htm

Add your comment:
Advertisement
Advertisement
Advertisement
Advertisement
Advertisement